# UAV Path Planning for Object Observation with Quality Constraints: A Dynamic Programming Approach

### v1

@@ Abstract
This paper addresses a UAV path planning task that seeks to observe a set of objects while satisfying the observation quality constraint. A dynamic programming algorithm is proposed that enables the UAV to observe the target objects with shortest path while subjecting to the observation quality constraint. The objects have their own facing direction and restricted observation range. With an observing order, the algorithm achieves $(1+\epsilon)$-approximation ratio in theory and runs in polynomial time. The extensive results show that the algorithm produces near-optimal solutions, the effectiveness of which is also tested and proved in the Airsim simulator, a realistic virtual environment.
Index Terms:
approximation, observation quality, path planning
I Introduction
@@
Benefiting from their flexibility and maneuverability in practice, UAVs generally require a trajectory to accomplish various tasks during flight. Coverage path planning algorithms are designed to determine feasible routes for UAVs to follow, covering a specific area or a set of Points of Interest (PoI) while satisfying certain conditions, such as distance constraints or hardware parameters, such as camera angles. Distance constraints refer to the UAV’s perception range, supported by cameras or radars, which require the target object to be located within a certain distance of the UAV. A wide range of algorithms has been proposed in the literature to enable UAVs to cover a specific space, planning feasible routes to cover PoIs (Points of Interest) or aggregating the area swept over during flight.
This study focuses on a realistic situation where a UAV must observe a set of objects while meeting a gross observation quality constraint, distinguishing it from a point cover problem. The problem can be extended to building inspection, where a path guides the UAV to capture photos of each side of the building and, meanwhile, considers the quality of the photos. However, manual planning of a route that is as short as possible can be challenging. To the best of our knowledge, our work is the first to propose a solution with observation quality and an approximation guarantee for UAV path planning, distinct from other literature that seldom considers the coverage quality and provides no theoretical bound on the path length. Our proposed solution improves the efficiency of UAV path planning, benefiting various industries and applications.
Related Work
As this paper plans a path for a UAV in covering problems, we present related works in three parts. The first part is devoted to different methods of UAV path planning. Then the focus is shifted to the UAV covering problem, in which the main task is to find a path that achieves covering effects. Finally, we briefly introduce several classical Traveling Salesman Problem (TSP) algorithms.
I-1 UAV path planning
@@
Sampling-based algorithms require prior knowledge of the map and divide it into a set of nodes. The concrete techniques used in these algorithms include Rapid-Exploring Random Trees (RRT), RRT-star, and A-star. Mathematical methods primarily employed Lyapunov functions [5] to maintain the stability of UAVs, linear programming to integrate all cost factors with Hamiltonian function to search for an optimal path, as described in [6], and Bezier curves, which further consider the smoothness of the flight path. Bio-inspired techniques [7] utilized evolutionary ideas by selecting a route as a parent path and generating new paths via mutation and crossover. The selection process is guided by the adaptive value of the offspring.
@@
AI-based methods have become a focal point of research in recent years, and these methods can be further classified into those based on traditional machine learning models, deep learning, and reinforcement learning. Supervised machine learning includes methods such as Gaussian filtering and Kalman filtering, which involve cleaning the data collected from the environment, inputting it into the model, and obtaining the predicted path and data visualization. Kang et al. in [8] and the authors in [9] used Kalman filters to clean noisy images collected by UAVs in complex environments, and obtained a safe flying space by outputting collision probabilities. However, due to Kalman filters’ limitations in recognizing numerous obstacles, the authors in  [10] used Gaussian filters to learn the impact of UAV maneuvers and more accurately estimate the UAV’s state. Unsupervised machine learning, such as the k-means method [11], determined the route by clustering the target points in a multi-task environment.
@@
Computer vision is a representative direction of deep learning applied to UAV path planning. Similar to the visual technology used in autonomous driving cars, UAVs use cameras to capture real-time images and identify possible obstacles and safe flying spaces [12][13]. In addition to deep learning, Q-learning in reinforcement learning was also used for path planning in [14], and Luan et al. in [15] used G-learning to iteratively optimize the path by calculating the cost function in real time. Lei et al. in [16] modeled the UAV navigation problem as a Markov decision process and proposed a model interpretation method based on feature attributes to explain the behavior of the UAV in the path planning process. Similarly, Xie in  [17] expressed path planning as a partially observable Markov process, constructing RNNs to tackle partially observable problems, and leveraging reward value and action value to reduce meaningless exploration. These two deep reinforcement learning methods significantly enhanced the learning efficiency of Q networks and G networks.
I-2 UAV covering
@@
As a subproblem of UAV path planning, UAV covering additionally requires achieving a cover task, as indicated in the name. One type is regional covering that assumes a region is covered when it falls within the perception area of the UAV. A relatively conventional algorithm is presented in [18] where Nam et al. utilized approximate cellular decomposition with criteria for both length and number of turns on the route. Yao et al.[19] proposed an offline route planning method based on Gaussian mixed model and heuristic prioritization, which was aimed at maximizing the probability of finding the lost target in a river rescue mission. Xie et al. [20] solved the problem of multiple polygon regions by integrating covering a single polygon and traveling salesman problem. These are the methods for one UAV and algorithms built on multiple UAVs are as follows. Jing et al.[1] navigated the UAV to inspect a large complex building structure by combining the set covering problem and the vehicle routing problem. Rapidly exploring random tree, as indicated by its name, is broadly used to explore an unknown region with high efficiency. In [21], a cooperative surveillance task was achieved by multiple UAVs where RRT was modified to find feasible trajectories passing suitable observing locations, on which particle swarm optimization was then performed. Focusing on the coverage path planning of heterogeneous UAVs, the authors in [22] established the UAV and regional model before using linear programming to accurately provide the best point-to-point flight path for each UAV. Then, inspired by the foraging behavior of ants, a heuristic algorithm based on ACS was proposed to search for the approximate optimal solution and to minimize the time consumption of tasks in the cooperative search system.
@@
The other type is object-oriented where the problem aims to cover a set of objects and the condition satisfied when the distance between the UAV and an object is less than a threshold. In this case, effective coverage can be treated as a binary in that 1 for covered and 0 for not covered, which can find applications in the context of wireless sensor network. Huang et al.[23] embedded turning angles and switching numbers during retrieving data into graph structure and obtained a path through the use of Generalized TSP solutions. Authors in [24] completed the data collection mission in a straight line situation, minimizing the UAV’s total flight time via dynamic programming and meanwhile retrieving a certain amount of data each sensor. After determining the possible flying path, Yang et al.[25] combined a genetic algorithm and ant colony optimization to select the optimal path for data collection. In [26], the agents viewed moving objects and obstacles as disks of different sizes and the goal was to find a collision-free coverage path in a dynamic changing environment while ensuring smoothness. As for multiple UAVs, Alejo et al. [27] integrated online RRT with genetic algorithms, guiding several UAVs collecting data simultaneously from sensors randomly distributed. A multi-agent architecture was designed in [28] for UAVs to patrol in a region and monitor key ground facilities.
I-3 Traveling Salesman Problem (TSP)
@@
    An Dynamic Programming based algorithm is designed which can solve the approximate shortest path given the quality constraint.
@@
    We design an Integer Linear Programming (ILP) which can compute the near-optimal solution via a solver with respect to epsilon.
@@
    We theoretically guarantee that the algorithm runs in polynomial time and achieves $1+\epsilon$ approximation given a visiting order.
@@
    We carry out various numerical and simulation experiments and the results show that our method produces near optimal solution and make significant improvement on the observation quality.
The structure of the rest of the paper is organized as follows. In Section II, the system model is presented followed by discretization and the formalization of the problem. The detail of our methods is elaborated on in Section III including the design of dynamic programming algorithm and determination of the observation orders. The process and results of the experiment are presented in Section IV.
II System Model and Problem Formulation
@@
II-A System Model
Consider a two-dimensional space in which there are $n$ objects and a UAV equipped with a camera that can move freely in this space. The objects set is referred to as $\mathcal{O}=\{o_{1},o_{2},o_{3}...o_{n}\}$. Each object $o_{i}$ faces a direction $\overrightarrow{d_{o_{i}}}$, the location of which is described by coordinates $(x_{o_{i}},y_{o_{i}})$. In order to ensure the observation quality, the UAV is not allowed to take photos from an over-deviated orientation, which is supported by the fact that the most information is obtained when the camera is facing directly against an object. Moreover, since distance also affects the observing quality, the UAV cannot observe the object if the distance exceeds a threshold. Concretely, the UAV has to be located within an angle range around the object’s direction in order to observe it, which is denoted as $\theta$. Besides, the maximum observation distance is defined to be $d_{max}$. To further ensure safety, the UAV cannot approach too close to the object, so the minimum distance is set to $d_{min}$. As illustrated in Figure. 1, the UAV can efficiently observes an object if and only if it satisfies Definition 1. Dashed blue lines indicate the effective area. We adapt the definition to our problem, which is based on another paper [36] in order to emphasize the quality of observation.
Definition 1 (Efficient observation).
@@ II-A System Model
We define the observation quality of the object $o_{i}$ as $q_{i}$, thus the gross quality of all objects $Q=\sum_{i}q_{i}$. There is also an observation quality constraint $q^{*}$ that the gross quality of all objects $Q$ has to attain. In formula, $Q\geq q^{*}$. Formula  (1) is presented to calculate $q_{i}$ drawn from  [36], which is short of $q(o_{i},p_{j})$, the observation quality of $o_{i}$ from the point $p_{j}$. Suppose that there are a maximum and minimum observation quality $q_{max}$, $q_{min}$ for a single object, which are decided by the observation area. To make the problem solvable, we introduce an additional assumption: $n*q_{min}\leq q^{*}\leq n*q_{max}$.
@@
Given that a single UAV is tasked with covering all specified objects from a starting point $P_{0}$, the objective is to minimize the path length while adhering to quality requirements. However, this poses a challenge as the problem cannot be solved directly through a Traveling Salesman Problem approach, which does not incorporate the unique characteristics of the observation points and quality. In certain cases, there may be an overlap in the observation areas of multiple objects, thereby enabling the UAV to capture images of multiple objects by rotating its camera at certain points. Due to the arbitrary nature of the object layout, devising a comprehensive solution that selects suitable observation points and plans the path remains a challenging task.
II-B Discretization of the observation points
Provided that the objects are located in a continuous space, a method is first designed to generate discrete observation points which enables the UAV to practically plan its flying path in the finite space. It is inevitable that discretization brings about $1+\epsilon$ approximation error [37]. We apply the method in [36] to our problem and it is briefly summarized below. Figure. 2 shows a possible discretization result of a target area.
It is intended to divide the whole effective observation area into several independent segments so that the observation quality in each segment is considered the same. Therefore, it only needs to choose the upper left corner point of a segment to represent it. The method begins by dividing the area in the distance domain from the object into $K1$ rings, with end points referred to as $l(0),l(1),...,l(K1)$. Each ring is further separated into $K2$ segments on both sides of $\overrightarrow{d_{o_{i}}}$, with endpoints in the angle domain defined as $a(0),...,a(K2)$. In Figure. 2, $K1$ is 3 and $K2$ equals 1,2,3, respectively for each sector ring. A point, for example, can be expressed by $(l(2),\alpha(1))$. With this method, it is ensured that the error of observation quality is associated with $\epsilon$. Apart from this, we make the following revision since the granularity may not be fine enough for the same approximation error of path length.
Let $A_{1}=l(k_{1}+1)-l(k_{1})$ refer to the length of a segment, and $B_{1}=\alpha(k_{2}+1)-\alpha(k_{2})$ refer to the angular width of the segment. Let $D$ denote the maximum distance between two objects from the set $\mathcal{O}$. Formally, $D=\max_{i,j\in|\mathcal{O}|}d\left(o_{i},o_{j}\right)$. We use Equation( 4) to determine the size of a segment and ensure the approximation ratio, with $n$ being the number of objects.
@@ II-B Discretization of the observation points
The two proposed equations are then used to generate all feasible segments and observation points in the space. The process is described in Algorithm 1. For convenience, we suppose there is an object $o_{0}$ centered at the original point and facing the positive x-axis direction ($\overrightarrow{d_{o_{0}}}=0$). The discretization method is then applied to $o_{0}$ that produces the observation point set $P_{*}$. For each object $o_{i}$, its corresponding observation points $P_{i}$ can be obtained by rotating all the points in $P_{*}$ to the direction of $o_{i}$ before translating to $(x_{o_{i}},y_{o_{i}})$. The aggregation of $P_{i}$ is $\mathcal{P}$ that contains $\{p_{1},\,p_{2},\,p_{3},\,...,\,p_{u}\}$ and the coordinate of the $i^{th}$ point is $(x_{p_{i}},\,y_{p_{i}})$. The observation points of an object are illustrated in Figure. 2, in which the orange dots are the observation points and the magnitude of observation quality can be indicated by the different blue color.
@@
Proof.
Given that observation points are utilized for approximating the points in a 2D continuous space, each point’s approximation error is limited to at most $\delta$ when rounding it to the nearest mesh point. Additionally, the total error is restricted to $n*\delta$, as no more than $n$ observation points are chosen in the algorithm, which will be elaborated later. It can be anticipated that the total error does not surpass $\epsilon$ times the lower bound of the optimal solution, ensuring that the actual relative error is at most $\epsilon$. As the UAV needs to travel at least this distance to observe two objects efficiently, and the situation with only one object is not considered, it is evident that $D$ fulfills the requirement. Thus, Lemma 1 is established. By setting $n*\delta=\epsilon*D$, Equation (4) is derived. ∎
II-C Integer Linear Programming representation

### v2

@@ Abstract
This paper addresses a UAV path planning task that seeks to observe a set of objects while satisfying the observation quality constraint. A dynamic programming algorithm is proposed that enables the UAV to observe the target objects with the shortest path while subjecting to the observation quality constraint. The objects have their own facing direction and restricted observation range. With an observing order, the algorithm achieves $(1+\epsilon)$-approximation ratio in theory and runs in polynomial time. The extensive results demonstrate that the algorithm produces near-optimal solutions, the effectiveness of which is also tested and proved in the Airsim simulator, a realistic virtual environment.
keywords:
multi-objective optimization , approximation algorithms , observation quality , path planning
journal: Computers and Electrical Engineering
[…]
[seu]organization=Southeast University, Department of Computer Science and Engineering, city=Nanjing, country=China
[…]
[ams]organization=Consulting Center for Strategic Assessment, Academy of Military Science, city=Beijing, country=China
1 Introduction
@@
Benefiting from their flexibility and maneuverability in practice, UAVs generally require a trajectory to accomplish various tasks during flight. Coverage path planning algorithms are specifically designed to determine feasible routes for UAVs to follow, covering either a specific area or a set of Points of Interest (PoI), while adhering to certain conditions, such as distance constraints or hardware parameters, such as camera angles. Distance constraints refer to the UAV’s perception range, supported by cameras or radars, which require the target object to be located within a certain distance of the UAV. A wide range of algorithms has been proposed in the literature to enable UAVs to cover a specific space, whether by planning feasible routes to cover PoIs (Points of Interest) or by aggregating the area swept over during flight.
This study delves into a realistic scenario where a UAV must observe a set of objects while meeting a gross observation quality constraint, distinguishing it from a point cover problem. The problem can be extended to building inspection, where a path guides the UAV to capture photos of each side of the building and, meanwhile, considers the quality of the photos. However, manual planning of a route that is as short as possible can be challenging. To the best of our knowledge, this work stands out as the first to propose a solution with observation quality and an approximation guarantee for UAV path planning, distinct from other literature that seldom considers the coverage quality and provides no theoretical bound on the path length. Our proposed solution improves the efficiency of UAV path planning, benefiting various industries and applications.
2 Related Work
As this paper plans a path for a UAV in covering problems, we present related works in three parts. The first part is devoted to different methods of UAV path planning. The focus is then shifted to the UAV covering problem, in which the main task is to find a path that achieves covering effects. Finally, we briefly introduce several classical Traveling Salesman Problem (TSP) algorithms.
2.1 UAV Path Planning
@@
Sampling-based algorithms require prior knowledge of the map and divide it into a set of nodes. The concrete techniques used in these algorithms include Rapid-Exploring Random Trees (RRT), RRT-star, and A-star. Mathematical methods primarily employed Lyapunov functions [5] to maintain the stability of UAVs, linear programming to integrate all cost factors with Hamiltonian function to search for an optimal path, as described in [6], and Bezier curves, which further consider the smoothness of the flight path. Bio-inspired techniques [7] utilized evolutionary ideas, selecting a route as a parent path and generating new paths through mutation and crossover. The selection process is guided by the adaptive value of the offspring.
@@
AI-based methods have become a focal point of research in recent years, and these methods can be further classified into those based on traditional machine learning models, deep learning, and reinforcement learning. Supervised machine learning includes methods such as Gaussian filtering and Kalman filtering, which involve cleaning the data collected from the environment, inputting it into the model, and obtaining the predicted path and data visualization. Kang et al. in [8] and the authors in [9] used Kalman filters to clean noisy images collected by UAVs in complex environments, and obtained a safe flying space by outputting collision probabilities. Due to Kalman filters’ limitations in recognizing numerous obstacles, however, the authors in  [10] used Gaussian filters to learn the impact of UAV maneuvers and more accurately estimate the UAV’s state. Unsupervised machine learning, such as the k-means method [11], determined the route by clustering the target points in a multi-task environment.
@@
Computer vision is a representative direction of deep learning applied to UAV path planning. Similar to the visual technology used in autonomous driving cars, UAVs use cameras to capture real-time images and identify possible obstacles and safe flying spaces [12],  [13]. In addition to deep learning, Q-learning in reinforcement learning was also used for path planning in [14], and Luan et al. in [15] used G-learning to iteratively optimize the path by calculating the cost function in real time. Lei et al. in [16] modeled the UAV navigation problem as a Markov decision process and proposed a model interpretation method based on feature attributes to explain the behavior of the UAV in the path planning process. Similarly, Xie in  [17] expressed path planning as a partially observable Markov process, constructing RNNs to tackle partially observable problems, and leveraging reward value and action value to reduce meaningless exploration. These two deep reinforcement learning methods significantly enhanced the learning efficiency of Q networks and G networks.
2.2 UAV Covering
@@
As a subproblem of UAV path planning, UAV covering additionally requires achieving a cover task, as indicated in the name. One type is regional covering that assumes a region is covered when it falls within the perception area of the UAV. A relatively conventional algorithm is presented in [18] where Nam et al. utilized approximate cellular decomposition with criteria for both length and number of turns on the route. Yao et al. [19] proposed an offline route planning method based on Gaussian mixed model and heuristic prioritization, which was aimed at maximizing the probability of finding the lost target in a river rescue mission. Xie et al. [20] solved the problem of multiple polygon regions by integrating covering a single polygon and traveling salesman problem. These are the methods for one UAV and algorithms built on multiple UAVs are as follows. Jing et al. [1] navigated the UAVs to inspect a large complex building structure by combining the set covering problem and the vehicle routing problem. Rapidly exploring random tree, as indicated by its name, is broadly used to explore an unknown region with high efficiency. In [21], a cooperative surveillance task was achieved by multiple UAVs where RRT was modified to find feasible trajectories passing suitable observing locations, on which particle swarm optimization was then performed. Focusing on the coverage path planning of heterogeneous UAVs, the authors in [22] established the UAVs and regional model before using linear programming to accurately provide the best point-to-point flight path for each UAV. Then, inspired by the foraging behavior of ants, a heuristic algorithm based on ACS was proposed to search for the approximate optimal solution and to minimize the time consumption of tasks in the cooperative search system.
@@
The other type is object-oriented where the problem aims to cover a set of objects, with the condition satisfied when the distance between the UAV and an object is less than a threshold. In this case, effective coverage can be treated as a binary in that 1 for covered and 0 for not covered, which can find many applications in the context of wireless sensor networks. Huang et al. [23] embedded turning angles and switching numbers during retrieving data into graph structure and obtained a path through the use of Generalized TSP solutions. Authors in [24] completed the data collection mission in a straight line situation, minimizing the UAV’s total flight time via dynamic programming and meanwhile retrieving a certain amount of data each sensor. After determining the possible flying path, Yang et al. [25] combined a genetic algorithm and ant colony optimization to select the optimal path for data collection. In [26], the agents viewed moving objects and obstacles as disks of different sizes and the goal was to find a collision-free coverage path in a dynamic changing environment while ensuring smoothness. Concerning multiple UAVs, Alejo et al. [27] integrated online RRT with genetic algorithms, guiding several UAVs collecting data simultaneously from sensors randomly distributed. A multi-agent architecture designed in [28] allowed UAVs to patrol in a region and monitor key ground facilities.
2.2.1 Traveling Salesman Problem (TSP)
@@
    A Dynamic Programming based algorithm is designed which can solve the approximate shortest path given the quality constraint.
@@
    An Integer Linear Programming (ILP) is designed which can compute the near-optimal solution via a solver with respect to epsilon.
@@
    It is theoretically guaranteed that the algorithm runs in polynomial time and achieves $1+\epsilon$ approximation given a visiting order.
@@
    We carry out various numerical and simulation experiments and the results show that our method produces near optimal solution and makes significant improvement on the observation quality.
The structure of the rest of the paper is organized as follows. In Section 3, the system model is presented followed by discretization and the formalization of the problem. The detail of our methods is elaborated on in Section 4 including the design of dynamic programming algorithm and determination of the observation orders. The process and results of the experiment are presented in Section 5.
3 System Model and Problem Formulation
@@
3.1 System Model
Consider a two-dimensional space in which there are $n$ objects and a UAV equipped with a camera that can move freely in this space. The objects set is referred to as $\mathcal{O}=\{o_{1},o_{2},o_{3},...,o_{n}\}$. Each object $o_{i}$ faces a direction $\overrightarrow{d_{o_{i}}}$, the location of which is described by coordinates $(x_{o_{i}},y_{o_{i}})$. In order to ensure the observation quality, the UAV is not allowed to take photos from an over-deviated orientation, which is supported by the fact that the most information is obtained when the camera is facing directly against an object. Moreover, since distance also affects the observing quality, the UAV cannot observe the object if the distance exceeds a threshold. Concretely, the UAV has to be located within an angle range around the object’s direction in order to observe it, which is denoted as $\theta$. Besides, the maximum observation distance is defined to be $d_{max}$. To further ensure safety, the UAV cannot approach too close to the object, so the minimum distance is set to $d_{min}$. As illustrated in Figure 1, the UAV can efficiently observe an object if and only if it satisfies Definition 1. Dashed blue lines indicate the effective area. We have adapted this definition [36] to underscore the emphasis on the quality of observation in our problem.
We define the observation quality of the object $o_{i}$ as $q_{i}$, thus the gross quality of all objects $Q=\sum_{i}q_{i}$. There is also an observation quality constraint $q^{*}$ that the gross quality of all objects $Q$ has to attain. In formula, $Q\geq q^{*}$. Formula (1) is presented to calculate $q_{i}$ drawn from  [36], which is short of $q(o_{i},p_{j})$, the observation quality of $o_{i}$ from the point $p_{j}$. Suppose that there exist maximum and minimum observation quality $q_{max}$, $q_{min}$ for a single object, which are decided by the observation area. To make the problem solvable, we introduce an additional assumption: $n*q_{min}\leq q^{*}\leq n*q_{max}$.
Definition 1 (Efficient observation)
@@ 3.1 System Model
We define the observation quality of the object $o_{i}$ as $q_{i}$, thus the gross quality of all objects $Q=\sum_{i}q_{i}$. There is also an observation quality constraint $q^{*}$ that the gross quality of all objects $Q$ has to attain. In formula, $Q\geq q^{*}$. Formula (1) is presented to calculate $q_{i}$ drawn from  [36], which is short of $q(o_{i},p_{j})$, the observation quality of $o_{i}$ from the point $p_{j}$. Suppose that there exist maximum and minimum observation quality $q_{max}$, $q_{min}$ for a single object, which are decided by the observation area. To make the problem solvable, we introduce an additional assumption: $n*q_{min}\leq q^{*}\leq n*q_{max}$.
@@
Given that a single UAV is tasked with covering all specified objects from a starting point $P_{0}$, the objective is to minimize the path length while adhering to quality requirements. However, this poses a challenge as the problem cannot be solved directly through a Traveling Salesman Problem approach, which does not account for the unique characteristics of the observation points and their associated qualities. In certain cases, there may be instances where observation areas of multiple objects overlap, allowing the UAV to capture images of multiple objects by rotating its camera at certain points. Due to the arbitrary nature of the object layout, devising a comprehensive solution that selects suitable observation points and plans the path remains a challenging task.
3.2 Discretization of the observation points
Provided that the objects are located in a continuous space, a method is first designed to generate discrete observation points which enables the UAV to practically plan its flying path in the finite space. It is inevitable that discretization brings about $(1+\epsilon)$ approximation error [37]. We apply the method in [36] to our problem and it is briefly summarized below. Figure 2 shows a possible discretization result of a target area.
It is intended to divide the whole effective observation area into several independent segments so that the observation quality in each segment is considered the same. Therefore, it only needs to choose the upper left corner point of a segment to represent it. The method begins by dividing the area in the distance domain from the object into $K_{1}$ rings, with end points referred to as $l(0),l(1),...,l(K_{1})$. Each ring is further separated into $K_{2}$ segments on both sides of $\overrightarrow{d_{o_{i}}}$, with endpoints in the angle domain defined as $a(0),...,a(K_{2})$. In Figure 2, there are $K_{1}=3$ rings and there are $K_{2}=\{1,2,3\}$ respectively for each sector ring. A point, for example, can be expressed by $(l(2),\alpha(1))$. With this method, it is ensured that the error of observation quality is associated with $\epsilon$. Apart from this, we make the following revision since the granularity may not be fine enough for the same approximation error of path length.
Let $A_{1}=l(k_{1}+1)-l(k_{1})$ refer to the length of a segment, and $B_{1}=\alpha(k_{2}+1)-\alpha(k_{2})$ refer to the angular width of the segment. Let $D$ denote the maximum distance between two objects from the set $\mathcal{O}$. Formally, $D=\max_{i,j\in|\mathcal{O}|}d\left(o_{i},o_{j}\right)$. We use Equation (4) to determine the size of a segment and ensure the approximation ratio, with $n$ being the number of objects.
@@ 3.2 Discretization of the observation points
The two proposed equations are then used to generate all feasible segments and observation points in the space. The process is described in Algorithm 1. For convenience, we suppose there is an object $o_{0}$ centered at the original point and facing the positive x-axis direction ($\overrightarrow{d_{o_{0}}}=0$). The discretization method is then applied to $o_{0}$ that produces the observation point set $P_{*}$. For each object $o_{i}$, its corresponding observation points $P_{i}$ can be obtained by rotating all the points in $P_{*}$ to the direction of $o_{i}$ before translating to $(x_{o_{i}},y_{o_{i}})$. The aggregation of $P_{i}$ is $\mathcal{P}$ that contains $\{p_{1},\,p_{2},\,p_{3},\,...,\,p_{u}\}$ and the coordinate of the $i^{th}$ point is $(x_{p_{i}},\,y_{p_{i}})$. The observation points of an object are illustrated in Figure 2, in which the orange dots are the observation points and the magnitude of observation quality can be indicated by the different blue color.
@@
Proof 1
Given that observation points are utilized for approximating the points in a 2D continuous space, each point’s approximation error is limited to at most $\delta$ when rounding it to the nearest mesh point. Additionally, the total error is restricted to $n*\delta$, as no more than $n$ observation points are chosen in the algorithm, which will be elaborated later. It can be anticipated that the total error does not surpass $\epsilon$ times the lower bound of the optimal solution, ensuring that the actual relative error is at most $\epsilon$. As the UAV needs to travel at least this distance to observe two objects efficiently, and the situation with only one object is not considered, it is evident that $D$ fulfills the requirement. Thus, Lemma 1 is established. By setting $n*\delta=\epsilon*D$, Equation (4) is derived.
3.3 Integer Linear Programming representation